OpenStudy (anonymous):

Evaluate the expression by using the given value of the variables. y ÷ 2 • 4 – 10 ÷ x + 6 where x = 5 and y = 16 i still cant get this one

OpenStudy (mathstudent55):

First step, replace the variables with the values you're given: \[16 \div 2 \times 4 - 10 \div 5 + 6\]

OpenStudy (mathstudent55):

Now you need to be careful with the order of operations. Multiplications and divisions have to be done before additions and subtractions.

OpenStudy (anonymous):

yes i did that but i keep getting 6 or 8

OpenStudy (mathstudent55):

So next step doing only multiplications and divisions you get: 32 - 2 + 6 Now do additions and subtractions 36

OpenStudy (anonymous):

i get it now thx for your help :)

OpenStudy (mathstudent55):

Multiplication and division have the same level of priority. Then addition and subtraction have the same level of priority.

OpenStudy (mathstudent55):

You'r welcome

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

follow the very basic rule every time called DMAS rule frist division then multiplication then addition and lastly substraction

OpenStudy (mathstudent55):

You can also call it PEMDAS and include parentheses and exponents. The order is: 1. Parentheses 2. Exponents 3. Multiplications and divisions in the order they appear from left to right 4. Additions and subtractions in the order they appear from left to right

OpenStudy (mathstudent55):

Look at: \[16 \div 2 \times 4\] 8 x 4 32

OpenStudy (mathstudent55):

The correct answer is 32 If you do multiplications before divisions you'd get: \[16 \div 2 \times 4\] \[16 \div 8\] 2

OpenStudy (mathstudent55):

In this case, 2 is an incorrect answer. 32 is the correct answer.

OpenStudy (goten77):

ignore what mariel said.. when it comes down to just adding/subtracting you can do it just across and get the same answer... no reason to do a extra step of organizing the adding then subtracting

OpenStudy (mathstudent55):

Goten77 also made a mistake when stating "m[ultiplication] comes before d[ivision]" That is not true. Multiplication and division are the same priority, so do them from left to right in the order they appear.

OpenStudy (goten77):

hmm i did didnt i

OpenStudy (mathstudent55):

Just be careful with it.

OpenStudy (goten77):

o i meant as in pemdas as in dont do like 2 of each at same time meaning dont do a (5x5/2x4) as in messing up by doing 5x5 all over 2x4....not sure how i worded that so badly

OpenStudy (anonymous):

16/2 x 4 - 10/5 + 6 8 x 4 - 2 + 6 32 - 2 + 6 30 + 6 36